### v3

@@ 5.4 Abatement Cost
$\mu_{i,t,j}^{U}=1-\frac{E_{i,t,j}^{U}}{E_{i,t,\mathrm{FFrun451}}^{U}},$ (28)
where $E_{i,t,\mathrm{FFrun451}}^{U}$ is the regional emission from the FFrun451 scenario with zero carbon tax (see Ueckerdt et al., (2019) for details). Then we use the computed $\mu_{i,t,j}^{U}$ and the associated carbon tax $\tau_{t,j}^{U}$ to find the abatement cost coefficients—$b_{1,i}$, $b_{2,i}$, $b_{3,i}$, and $b_{4,i}$such that the equality (27) can hold in an approximate manner for every scenario $j$ and time $t$.
6 Results
In this section, we discuss our simulation results. We first present the results of the noncooperative model under the baseline emission cap scenario for the period from 2020 to 2100. Next, we compare the economic outcomes across different models, examining the role of the emission caps and the implementation of a global ETS. Lastly, we present simulation results with the different emission cap scenarios assuming the net zero target year to be 2050, 2070, and 2090, respectively, for all 12 regions.
6.1 Key Results under the Baseline Emission Caps
@@
Figure 1 displays simulation results at the global level, from the noncooperative model with the ETS under the baseline emission cap scenario. Until 2100, we can delineate three periods according to the global emission trading patterns. Prior to 2025, global emissions do not reach the global emission caps and therefore the permit price remains zero, implying an excess supply of emission permits in the initial years (Figure 1 top-left and bottom-left), as the emission caps in the initial years are large while they become much smaller over time (Figure A.5). This result is not surprising, considering that an oversupply of emission permits has been observed in the EU ETS (Fuss et al.,, 2018), resulting in zero or very low permit prices. This is also consistent with our theoretical analysis for the regional social planners’ Nash equilirbium in Section 2. During this period, the volume of global emission abatement and its associated abatement cost remain at low levels (Figure 1 top-left and bottom-right). This result implies that the global emission cap should be set at a level such that there is no over-supply of emission permits, so that permit prices are strictly positive under noncooperation and the ETS.
@@
Global emissions are constrained by the global emission caps starting from 2025. To comply with the monotonically decreasing emission caps, the regions undertake additional abatement efforts and/or purchase emission permits. This results in a substantial increase in the volume of emission abatement, along with the abatement cost and permit price. The decrease in global emissions is mainly achieved by the concomitant increase in global abatement, which peaks by around 2060 (Figure 1 top-left). The global abatement cost also increases steeply to reach its maximum value of \$5.45 trillion by 2070, and then decreases as the world reaches net zero emissions by 2070. The emission permit trade gradually decreases to zero by 2070 (Figure 1 top-left). As the emission cap becomes tighter over the years, the permit price increases to \$845 per ton of carbon in 2050 and \$2,059 in 2069 (Figure 1 bottom-left). The kinks in the permit price path in 2050 and 2060 are a result of some regions achieving net zero emissions, as shown in Figure A.5. Essentially, binding emission caps lead to a rise in the overall emission abatement, while the steep increase in the permit price limits the trading of emission permits.
@@
Post-2070 is the period when global emissions are net zero. For this period, emission permits are no longer traded and all emissions are abated in each region. [^16] In the bottom-left panel of Figure 1, we plot the emission permit price only when the traded volume is positive. By the end of this century, the atmospheric temperature is projected to reach 1.7°C above the pre-industrial level, which is driven by net positive emissions before 2070 (Figure 1 top-right), restricted by the emission caps. [^17] If there is no emission cap, then the temperature anomaly will be much higher. This can be seen in a later discussion with alternative emission caps. Lastly, the global climate damage rises almost linearly over the entire period (Figure 1 bottom-right). Overall, our noncooperative model simulation predicts that the global emission caps set by the Paris Agreement and Glasgow Pact are not restrictive enough to achieve the global target of limiting the temperature rise to 1.5°C.
It is instructive to examine the trading patterns of different regions in the emission permit market over the years to identify the permit buyers and sellers. Figure 2 displays the volumes of traded emission permits for each region over time, with a positive value denoting permit purchase and a negative value indicating permit sales. Before 2025, although there is an excess permit supply at the global level, the emission cap constraint is effective for some regions, such as China and Latin America, as these regions become permit buyers at this early stage. After 2025, when the global emission cap constraint becomes binding, the group of permit buyers consists of the US, EU, India, and OHI regions; the group of permit sellers consists of Russia, China, and Eurasia, with China being the largest permit provider after 2036. Japan is expected to be involved in a relatively small volume of permit trading, and the other regions (Africa, MidEast, Latin America, and OthAs) change their status of permit suppliers to buyers or vice versa over time.
@@
6.2 SCC, MAC, and the Optimal Carbon Tax
Based on the simulation results under the baseline emission cap scenario, this section examines the relationship between the MAC, regional social cost of carbon (SCC), and the optimal carbon tax under the global ETS regime. The MAC captures the additional cost incurred due to an increase in emission abatement. From our model equation (10), the total abatement cost is $\mathrm{TAC}{}_{i,t}=b_{1,i,t}\mu_{i,t}^{b_{2,i}}Q_{i,t},$ in trillions of USD. Thus, the MAC, in 2020 USD per ton of carbon, is obtained as follows:
$\mathrm{MAC}{}_{i,t}=1,000\left(\frac{\partial\mathrm{TAC}{}_{i,t}}{\partial E_{i,t}^{A}}\right)=1,000\left(\frac{b_{1,i,t}b_{2,i}\mu_{i,t}^{b_{2,i}-1}}{\sigma_{i,t}}\right).$ (29)
With $b_{1,i,t}>0$ and $b_{2,i}>2,$ MAC is strictly increasing on emission control rate $\mu_{i,t}\in[0,1].$ [^18] See Table A.2 in Appendix A.1 for the list of our calibrated abatement cost parameters. Figure 3 displays the regional MAC along with the market equilibrium price of emission permits. We find that, for every region, the MAC increases rapidly and remains strictly greater than the permit price until the emissions hit zero by 2050s or 2060s. [^19] For regional optimal emissions, see Figure 11 in Appendix A.4.1. As discussed in Section 2, $\mathrm{MAC}(E_{i,t}^{*})>m_{t}$ implies that the optimal emissions for the regional social planner do not align with the optimal emissions for the representative firm under no tax in each region. Therefore, the optimal regional carbon tax can be introduced at $\tau_{i,t}=\mathrm{MAC}(E_{i,t}^{*})-m_{t}$, so that the firm’s emissions are equal to the regional social planner’s optimal emissions $E_{i,t}^{*}$, for region $i$ at time $t$ before 2050 or 2070. Russia is an exception, where the MAC falls below the permit price starting in 2025, even though it has nonzero emissions until 2050, because global warming leads to increase in productivity rather than causing climate damages in Russia. This result is strongly influenced by the climate damage parameters, calibrated using projections from Burke et al., (2018).
@@ 6.2 SCC, MAC, and the Optimal Carbon Tax
In Figure 3, the MAC of each region gradually decreases below the permit price after its net zero emissions is achieved (i.e., $E_{i,t}^{*}=0$). Russia is the first region to achieve net zero emissions in 2050 (as shown in Figure A.7 in Appendix A.4.1). However, Russia continues to sell permits afterwards, as shown in Figure 2, because its net zero emission target year is 2060, under the baseline emission cap scenario. After 2050, Russia’s MAC begins to decline slightly and stabilizes over time, as its emission control rate reaches its upper limit. [^20] When emission control rates are one, the MACs are $1,000b_{2,i}(b_{1,i}+b_{3,i}\exp(-b_{4,i}t))$, and they are nearly constant when $t$ is large. Similar trends are followed by other regions as they attain net zero emissions in the 2050s ( China, the US, the EU, the OHI, MidEast, Eurasia, and Latin America) and the 2060s (Japan and India). Africa and OthAs are the last group of regions to achieve net zero emissions (before trading of permits) by 2070, after which emission permits are no longer traded, and the MACs of all regions decline slowly over time. Recall that, when the regional net zero emissions (before trading of permits) are achieved (i.e., $E_{i,t}^{*}=0$), the regional optimal carbon tax should be larger than the gap between the regional MAC and permit price, i.e., $\tau_{i,t}\geq\mathrm{MAC}(E_{i,t}^{*})-m_{t}$. Note that if a region’s after-permit-trade zero emission cap constraint is binding (i.e., the inequality (12) is binding with $\overline{E}_{i,t}=0$), then even when its regional MAC is smaller than the permit price, the region will not be able to sell emission permits, otherwise it will violate the constraint.
Given the simulation results for the optimal carbon tax, we now turn to discuss the relationship between the optimal carbon tax and the regional social cost of carbon (SCC). The SCC, a central concept in the climate change literature, is widely used to quantify the present value of climate damages induced by an additional unit of carbon emissions. While the SCC is often calculated in a global social planner’s problem (e.g., the DICE model), we consider the SCC for each region. Similar to van der Ploeg and de Zeeuw, (2016) and Cai et al., (2023), we define the noncooperative SCC of a region as the marginal rate of substitution between global emissions and regional capital as follows:
@@
for each region $i$ under the open loop Nash equilibrium. Since our cumulative global emissions are measured in gigatonnes of carbon (GtC) and capital is measured in trillions of USD, our SCC is measured in 2020 USD per ton of carbon. [^21] To compute the regional SCC in the noncooperative model, it is equivalent to replace the numerator in equation (30) with the shadow price of the transition equation of cumulative global emissions (14) at time $t$, and replace the denominator with the shadow price of the regional capital transition equation (20) for each region.
Our concept of the regional SCC differs from the definition in Nordhaus, (2017) and Ricke et al., (2018), where the regional SCC is defined as the present value of future climate damages in a region resulting from an additional unit of global emissions released in the current period, but their social discount rates are exogenous. We define the regional noncooperative SCC as the marginal rate of substitution between cumulative global emissions and regional capital under the Nash equilibrium. With our definition, the regional noncooperative SCC is equal to the optimal regional carbon tax under noncooperation when the regional emissions are not binding at zero (i.e., the regional emission control rate is less than one), or is larger than the optimal regional carbon tax when the regional emissions are binding at zero. This pattern is consistent with the relation between the global SCC and the optimal global carbon tax as shown in Cai et al., (2017) and Cai and Lontzek, (2019).  [^22] We also verified numerically that our regional noncooperative SCC is equal to the present value of future climate damages in the region resulting from an additional unit of global emissions released in the current period, with our social discount rates $r_{i,t+1}$ defined endogenously with the following formula: $r_{i,t+1}=\frac{u^{\prime}(c_{i,t})}{\beta u^{\prime}(c_{i,t+1})}-1,$ where $c_{i,t}$ are the optimal per-capita consumptions.
@@
The simulation results confirm the relationship between the optimal regional carbon tax and the regional SCC under noncooperation with the ETS. Figure 4 shows that when the emissions of a region are strictly positive, the regional SCC is exactly equal to the optimal carbon tax, which is equal to the gap between the regional MAC and the permit price in Figure 3. For example, in 2050, the MAC for the US is \$1,081 per ton of carbon, the permit price is \$845, and their difference is exactly equal to the regional SCC of the US, \$236. Figure 4 also shows that when the regional emissions are zero, the regional SCC is larger than the gap, which could be negative. Among all the regions, the US has the highest SCC in the near term, at around \$198 per ton of carbon in 2030. However, the SCCs of Africa and non-OECD Asia are expected to experience substantial increases, reaching \$682 for Africa and \$610 for non-OECD Asia by the end of the century. Russia, with the lowest SCC, experiences a steady negative SCC, reaching (-\$88) per ton of carbon by the end of the century, indicating that it benefits from global warming throughout the entire period. Our results show considerable heterogeneity in the SCC across regions and suggest that significant regional carbon tax implementation is required, not only for developed countries such as the US, but also for developing countries like Africa, India, and non-OECD Asia under the global ETS with the baseline emission cap scenario.
@@ 6.3 Effects of ETS Implementation
Next, we examine the economic and climate implications of the ETS implementation by comparing the global economy with and without the ETS regime, while maintaining the baseline emission caps in both economies. The top two panels of Figure 5 show that the ETS implementation results in slightly higher emissions and temperature increases compared to the case without the ETS. This occurs because, under the ETS, regions with binding regional emission cap constraints now have the option to purchase permits from regions with less restrictive emission caps, fully exploiting the total amount of permits allowed at the global level, but the buyers do not have this option under no ETS. That is, the global emissions could be less than the global emission caps under no ETS, while they are equal under the ETS. This pattern persists until 2043. In the bottom two panels of Figure 5, we compare the MAC and the SCC using the US as an illustrative example. The MAC of the US economy is higher without the ETS until 2057, reflecting that the US becomes a permit buyer under the ETS regime as MAC increases with increasing emission abatement. The SCC comparison demonstrates that the ETS has little impact on the regional SCC. This pattern holds for the other regions as well, as shown in Figure A.10. Since the regional noncooperative SCC is the present value of future climate damages in the region resulting from an additional unit of global emissions released in the current period, these results show that the ETS has little impact on the marginal damages and our endogenous social discount rates.
@@
What are the welfare implications of implementing the global ETS? To quantify these effects, we compute a compensating variation (CV) per capita associated with the ETS implementation. Specifically, the CV per capita for region $i$ is computed by solving the following equation:
@@
is the social welfare associated with the vector of consumption per capita $\mathbf{c}_{i}=(c_{i,0},\cdots,c_{i,t},\cdots,c_{i,T})$ with the terminal time $T=299$. Note that a direct comparison of the noncooperative model with the ETS to the noncooperative model without the ETS does not provide a clear picture of the welfare effects of the ETS under the current baseline emission cap scenario, as the global emissions and resulting climate damages differ across the two economies. That is, the differences between $\mathbf{c}_{i}^{1}$ and $\mathbf{c}_{i}^{0}$ depend on both the emission caps and the ETS implementation. Compared with the economy without the ETS, temperature under the ETS implementation is expected to be higher as shown in Figure 5. Consequently, some regions may experience additional climate damage from the difference of temperatures between two cases, while it is greater than the reduced abatement cost from purchasing permits or the additional profit from selling permits under the ETS. This leads to negative welfare effects of the ETS implementation for some regions. To isolate the impact of the ETS, we adjust the emission caps $\overline{E}_{i,t}$ in the noncooperative model with the ETS to $\overline{E}^{\prime}_{i,t}$, which are the optimal emission levels obtained from the noncooperative model with the emission caps $\overline{E}_{i,t}$ but without the ETS, then we compare economies with and without the ETS under the emission caps $\overline{E}^{\prime}_{i,t}$ as these two economies have the same global emission levels and temperatures.
Table 1 shows the CV per capita in 2020 USD, and its share (%) out of per capita consumption in each region at $t=0$, when the ETS implementation does not impact temperature. The findings indicate that the welfare impacts of the global ETS are strictly positive for all regions, irrespective of whether they are permit sellers or buyers, with considerable heterogeneity in the magnitude of these effects across the regions. For example, in the United States, the CV per capita is \$87.22, which is a relatively small fraction of per capita consumption (0.2%). Among the 12 regions, Russia experiences the largest welfare improvement, with a CV per capita of \$183.53, equivalent to 2.2% of its per capita consumption. In contrast, Africa and non-OECD Asia have the smallest welfare gains from the global ETS. These results highlight that the principle of gains from trade applies to emissions trading as well, driven by efficiency gains achieved through the reallocation of emission abatement efforts across regions.
Table 1: Welfare effects of the ETS implementation.

### v4

@@ 6.4 Abatement Cost
$\mu_{i,t,j}^{U}=1-\frac{E_{i,t,j}^{U}}{E_{i,t,\mathrm{FFrun111}}^{U}},$ (21)
[…]
6.5 Climate System: Transient Climate Response to Emissions
In the TCRE climate system, $\mathcal{E}_{0}$, is chosen such that the initial global mean temperature $T_{0}$ in year 2020 is 1.2 degrees Celsius above the pre-industrial level. The contribution rate of cumulative global emissions to temperature is calibrated at $\zeta=0.0021$ by using the projections of emissions and temperatures across the four Representative Carbon Concentration Pathways (RCPs) scenarios: RCP 2.6, RCP 4.5, RCP 6.0, and RCP 8.5 (Meinshausen et al.,, 2011). That is, we solve the following minimization problem:
$\min_{\zeta}$ $\sum_{t}\sum_{j}\left\|T_{t}^{j}-\zeta\mathcal{E}_{t}^{j}\right\|$
$s.t.$ $\mathcal{E}_{t+1}^{j}=\mathcal{E}_{t}^{j}+E_{t}^{j},\ \forall t,j,$
where $j$ represents one of four RCP scenarios, $E_{t}^{j}$ and $T_{t}^{j}$ are the exogenous projections of emissions and temperatures, respectively, at time $t$ for RCP scenario $j$. Figure A.1 in Appendix A.3.1 demonstrates that the calibrated TCRE climate system matches well all four RCP temperature pathways using their associated RCP emission pathways.
7 Baseline Model Results
This section presents the numerical simulation results under the assumption that all regions participate in the global ETS. The baseline emission cap scenario is defined in accordance with the NDCs in the near term and the net-zero targets in the mid- to long term. We report the results through the end of the century (2020–2100), focusing on both economic and environmental outcomes across key variables. Additionally, we provide an in-depth analysis of the numerical simulation results to evaluate the effects of implementing a global cap-and-trade system.
7.1 Emissions, Temperature, and Permit Prices
@@
Figure 1 displays key simulation results at the global level, from the noncooperative model with the ETS under the baseline emission cap scenario. Until 2100, we can delineate three periods according to the global emission trading patterns. Prior to 2024, global emissions do not reach the global emission caps and therefore the permit price remains zero, implying an excess supply of emission permits in the initial years (Figure 1 top-left and bottom-left), as the emission caps in the initial years are large while they become much smaller over time (Figure A.5). This result is not surprising, considering that an oversupply of emission permits has been observed in the EU ETS (Fuss et al.,, 2018), resulting in zero or very low permit prices. During this period, the volume of global emission abatement and its associated abatement cost remain at low levels (Figure 1 top-left and bottom-right). This result implies that the global emission cap should be set at a level such that there is no over-supply of emission permits, so that permit prices are strictly positive under noncooperation and the ETS.
@@
Global emissions are constrained by the global emission caps starting from 2024. To comply with the monotonically decreasing emission caps, the regions undertake additional abatement efforts and/or purchase emission permits. This results in a substantial increase in the volume of emission abatement, along with the abatement cost and permit price. The decrease in global emissions is mainly achieved by the concomitant increase in global abatement, which peaks by around 2070 (Figure 1 top-left). The global abatement cost also increases steeply to reach its maximum value of \$7.82 trillion by 2070, and then decreases as the world reaches net zero emissions by 2070. The emission permit trade also gradually decreases to zero by 2070 (Figure 1 top-left). As the emission cap becomes tighter over the years, the permit price increases to \$923 per ton of carbon in 2050 and \$2,105 in 2069 (Figure 1 bottom-left). The kinks in the permit price path in 2030 result from a kink in emission cap pathways (see Figure A.6), while those in 2050 and 2060 are a result of some regions achieving net zero emissions, as shown in Figure A.5. Essentially, binding emission caps lead to a rise in the overall emission abatement, while the steep increase in the permit price limits the trading of emission permits.
@@
Post-2070 is the period when global emissions are at net zero. For this period, emission permits are no longer traded and all emissions are abated in each region. [^19] In the bottom-left panel of Figure 1, we plot the emission permit price only when the traded volume is positive. By the end of this century, the atmospheric temperature is projected to reach 1.7 degrees Celsius above the pre-industrial level, which is driven by net positive emissions before 2070 (Figure 1 top-right), restricted by the emission caps. [^20] If there is no emission cap, then the temperature anomaly will be much higher. This can be seen in a later discussion with alternative emission caps. Lastly, the global climate damage rises almost linearly over the entire period (Figure 1 bottom-right). Overall, our noncooperative model simulation predicts that the global emission caps set by the Paris Agreement and Glasgow Pact are not restrictive enough to achieve the global target of limiting the temperature rise to 1.5 degrees Celsius.
It is instructive to examine the trading patterns of different regions in the emission permit market over the years to identify the permit buyers and sellers. Figure 2 displays the volumes of traded emission permits for each region over time, with a positive value denoting permit purchase and a negative value indicating permit sales. Before 2024, although there is an excess permit supply at the global level, the emission cap constraint is effective for some regions, such as China and Latin America, as these regions become permit buyers at this early stage. After 2024, when the global emission cap constraint becomes binding, the group of permit buyers consists of the US, EU, India, and OHI regions; the group of permit sellers consists of Russia, China, and Eurasia, with China being the largest permit provider after 2032. Japan is expected to be involved in a relatively small volume of permit trading, and the other regions (Africa, MidEast, Latin America, and OthAs) change their status of permit suppliers to buyers or vice versa over time.
@@
7.2 Regional MAC and SCC
Based on the simulation results under the baseline emission cap scenario, this section examines the relationship between the regional MAC, SCC, and the market equilibrium emissions permit price under the global ETS regime. The MAC captures the additional cost incurred due to an increase in emission abatement. From our model equation (8), the total abatement cost is $\Phi_{i,t}=b_{1,i,t}\mu_{i,t}^{b_{2,i}}Q_{i,t},$ in trillions of USD. Thus, the MAC, in 2020 USD per ton of carbon, is obtained as follows:
$\mathrm{MAC}{}_{i,t}=1,000\left(\frac{\partial\Phi_{i,t}}{\partial E_{i,t}^{A}}\right)=1,000\left(\frac{b_{1,i,t}b_{2,i}\mu_{i,t}^{b_{2,i}-1}}{\sigma_{i,t}}\right).$ (22)
Note that, with $b_{1,i,t}>0$ and $b_{2,i}>2,$ MAC is strictly increasing on emission control rate $\mu_{i,t}\in[0,1].$ [^21] See Table A.2 in Appendix A.1 for the list of our calibrated abatement cost parameters. Figure 3 displays the regional MAC along with the market equilibrium price of emission permits. The simulation result shows that, for most regions, the MAC increases rapidly and remains strictly greater than the permit price until the emissions hit zero by 2050s or 2060s.  [^22] For regional optimal emissions, see Figure A.7 in Appendix A.4.1. Russia is an exception, where the MAC falls below the permit price starting in 2021, which we will elaborate shortly in the next paragraph. The MAC of each region gradually decreases below the permit price after its net zero emission level is achieved (i.e., $E_{i,t}^{*}=0$). Russia is the first region to achieve net zero emissions in 2048 (as shown in Figure A.7 in Appendix A.4.1). However, Russia continues to sell permits afterwards, as shown in Figure 2, because its net zero emission target year is 2060, under the baseline emission cap scenario. After 2048, Russia’s MAC begins to decline slightly and stabilizes over time, as its emission control rate reaches its upper limit. [^23] When emission control rates are one, the MACs are $1,000b_{2,i}(b_{1,i}+b_{3,i}\exp(-b_{4,i}t))$, and they are nearly constant when $t$ is large. Similar trends are observed in other regions as they attain net zero emissions in the 2050s (China, Latin America, MidEast, the US, the OHI, Eurasia) and the 2060s (the EU, Japan and India). Africa and OthAs are the last group of regions to achieve net zero emissions (before trading of permits) by 2070, after which emission permits are no longer traded, and the MACs of all regions decline slowly over time. Note that if a region’s after-permit-trade zero emission cap constraint is binding (i.e., the inequality (10) is binding with $\overline{E}_{i,t}=0$), then even when its regional MAC is smaller than the permit price, the region will not be able to sell emission permits, otherwise it will violate the constraint.
@@ 7.2 Regional MAC and SCC
The SCC, a central concept in the climate change literature, is widely used to quantify the present value of climate damages induced by an additional unit of carbon emissions. While the SCC is often calculated in a global social planner’s problem (e.g., the DICE model), we consider the SCC for each region. Similar to van der Ploeg and de Zeeuw, (2016) and Cai et al., (2023), we define the noncooperative SCC of a region as the marginal rate of substitution between global emissions and regional capital as follows:
@@
for each region $i$ under the open loop Nash equilibrium. [^24] To compute the regional SCC in the noncooperative model, it is equivalent to replace the numerator in equation (23) with the shadow price of the transition equation of cumulative global emissions (12) at time $t$, and replace the denominator with the shadow price of the regional capital transition equation (5) for each region. Since our cumulative global emissions are measured in gigatonnes of carbon (GtC) and capital is measured in trillions of USD, our SCC is measured in monetary unit of 2020 USD per ton of carbon.  [^25] The choice of discount rates has been a critical factor contributing to the gap in social cost of carbon estimates across different studies (Guo et al.,, 2006; Weitzman,, 2013). Our concept of regional SCC differs from the definition in Nordhaus, (2017), where the regional SCC is obtained by assuming that it is a fraction of the global SCC, with the share determined by the discounted value of regional output using an exogenous, constant discount rate. We verified numerically that our regional noncooperative SCC is equal to the present value of future climate damages in the region resulting from an additional unit of global emissions released in the current period, with our social discount rates $r_{i,t+1}$ defined endogenously with the following formula: $r_{i,t+1}=\frac{u^{\prime}(c_{i,t})}{\beta u^{\prime}(c_{i,t+1})}-1,$ where $c_{i,t}$ are the optimal per-capita consumptions.
The simulation results confirm the relationship between the regional SCC, MAC, and permit prices in a multi-region economy with a global ETS, as explored in Section 3. Figure 4 demonstrates that, when a region’s net emissions are strictly positive, the regional SCC is exactly equal to the gap between the regional MAC and the permit price shown in Figure 3.262626When regional net emissions are at zero, the regional SCC can be larger than the gap, $\left(\mathrm{MAC}_{i}(E_{i})-m^{*}\right)$, which could be negative, as shown in Figure 4. For example, in 2050, the MAC for the US is \$1,159 per ton of carbon, the permit price is \$923, and their difference is exactly equal to the regional SCC of the US, \$236. Russia is an exceptional case with a negative SCC in our simulation, suggesting that global warming creates benefits rather than causing climate damages in Russia, a result heavily influenced by the climate damage parameters calibrated using projections from Burke et al., (2018). Nevertheless, the relationship between the SCC, MAC, and permit price remains intact; for example, in 2040, the SCC for Russia is $\mathdollar 41$, which corresponds to the difference between the MAC (\$602) and the permit price (\$643). This explains why the MAC falls below the permit price starting in 2021, despite Russia having nonzero emissions until 2050. Among all the regions, the US has the highest SCC in the near term, at around \$199 per ton of carbon in 2030. However, the SCCs of Africa and non-OECD Asia are expected to experience substantial increases, reaching \$675 for Africa and \$599 for non-OECD Asia by the end of the century. Russia, with the lowest SCC, experiences a steady negative SCC, reaching (-\$88) per ton of carbon by the end of the century, indicating that it benefits from global warming throughout the entire period. Our results show considerable heterogeneity in the SCC across regions.
@@ 7.3 Effects of ETS Implementation
Next, we examine the economic and climate implications of the ETS implementation by comparing the global economy with and without the ETS regime, while maintaining the baseline emission caps in both economies. The top two panels of Figure 5 show that the ETS implementation results in slightly higher emissions and temperature increases compared to the case without the ETS. This pattern persists until 2043. This occurs because, under the ETS, regions with binding regional emission cap constraints now have the option to purchase permits from regions with less restrictive emission caps, fully exploiting the total amount of permits allowed at the global level. In contrast, without the ETS, regions with binding regional emission caps cannot utilize surplus emission permits from other regions, resulting in global net emissions that are lower than or equal to those in the ETS scenario over time. In the bottom two panels of Figure 5, we compare the MAC and the SCC using the US as an illustrative example. The MAC of the US economy without the ETS is higher until 2056, reflecting that the US becomes a permit buyer under the ETS regime as MAC increases with increasing emission abatement. [^27] Before 2030, the emission control rate grows fast so the MAC of the US increases rapidly along time, but after 2030, the emission control rate grows slowly such that the improvement of emission abatement technology makes the MAC decline along time until 2042. The SCC comparison demonstrates that the ETS has little impact on the regional SCC. This pattern holds for the other regions as well, as shown in Figure A.10. Since the regional noncooperative SCC is the present value of future climate damages in the region resulting from an additional unit of global emissions released in the current period, these results indicate that the ETS has little impact on either the marginal climate damages or our endogenous social discount rates.
@@
What are the welfare implications of implementing the global ETS? To quantify these effects, we compute a compensating variation (CV) per capita associated with the ETS implementation. Specifically, the CV per capita for region $i$ is computed by numerically solving the following equation:
@@
is the social welfare associated with the vector of consumption per capita $\mathbf{c}_{i}=(c_{i,0},\cdots,c_{i,t},\cdots,c_{i,T})$ with the terminal time $T=299$.
Table 1 shows welfare effects of the global ETS, measured by the CV per capita in 2020 USD, and its share (%) out of per capita consumption in each region at $t=0$. The results reveal significant heterogeneity in welfare effects across regions: some regions experience welfare gains, with Russia benefiting the most, showing a CV per capita of \$552.75, equivalent to 6.699% of its per capita consumption. In contrast, other regions, including Africa, the Middle East, and non-OECD Asia, experience negative welfare effects. This outcome may seem counterintuitive, as one might expect an additional market mechanism (ETS) to enhance welfare by providing greater flexibility in managing emissions. It is important to note that a direct comparison of the noncooperative model with and without the ETS does not provide a clear picture of the welfare effects of the ETS under the current baseline emission cap scenario, as the global net emissions and resulting climate damages differ across the two economies. Recall that, global average temperature is higher under ETS implementation in the baseline emission cap scenario, as shown in Figure 5. Therefore, the differences between $\mathbf{c}_{i}^{1}$ and $\mathbf{c}_{i}^{0}$ depend on both the global net emissions (and resulting climate damages) and the ETS implementation. As a result, some regions may experience additional climate damages due to higher temperatures, outweighing the benefits of lower abatement costs from permit purchases or additional profits from permit sales. This explains why some regions experience negative welfare effects from ETS implementation.
Table 1: Welfare effects of the ETS implementation: pre-emission cap adjustment.